OpenStudy (summersnow8):

#38 A gasoline station in Manila, Philippines, charges 38.46 pesos per liter of unleaded gasoline at a time when one U.S dollar (USD) buys 47.15 pesos (PHP). The car you are driving has a gas tank capacity of 14 U.S. gallons and gets 24 miles per gallon. How do you solve these? Answer to A is USD 3.09 Answer to B is USD 43 Answer to C is 2.1 x 102 mi

OpenStudy (summersnow8):

A. What is the cost of unleaded gasoline in Manila in USD per gallon? B. How much would a tankful of unleaded gasoline for your car cost in USD? C. Suppose that you have only PHP 1255 (a day’s wage for an elementary school teacher) and the car’s tank is almost empty. How many miles can you expect to drive if you spend all your money on gasoline?

OpenStudy (anonymous):

What do you think needs to be done to answer A? In C did you mean 12.55 PHP?

OpenStudy (cuanchi):

The answers are correct. I dont know if you have to express your answers with the correct number of significant figures, A has to have 4SF and B and C 2 significant figures

OpenStudy (summersnow8):

@Cuanchi I know the answers are correct, because they are in the back of the book.... I need to understand how to get those answers though!

OpenStudy (summersnow8):

@LeibyStrauss I am not sure, but for C, the book says "PHP 1255"

OpenStudy (anonymous):

I am working them one by one. Once I figured it out all out, I will post it. Do not log off though!

OpenStudy (summersnow8):

@Shalante Thank you!

OpenStudy (anonymous):

1. You always have to start with given units and try to solve for unknown units by canceling out the given & let the unknown units remain. \[\frac{ peso }{ liter}\times \frac{ USD }{ peso }\times \frac{ liter }{ gallon }\] If lets say 1USD=38.46 pesos You can put it conversion in both ways like this \[\frac{ 38.6 pesos }{ 1USD } or \frac{ 1USD }{ 38.6pesos }\]

OpenStudy (anonymous):